Ask
your own question, for FREE!
Mathematics
9 Online
help? http://prntscr.com/9et7n7
Still Need Help?
Join the QuestionCove community and study together with friends!
@Directrix
@Tommynaiter
1. The probability of picking a red marble is 5/15 = 1/3: desired/total. If you replace the marble, the probability of picking another red marble is the same, 5/15 = 1/3. The probability of picking red then red is 1/3 * 1/3 = 1/9. 2. The probability of picking the first red marble is 3/10. Since you don't replace the marble, the probability of picking the second red marble is 2/9. The probability of picking red then red is 3/10 * 2/9 = 6/90 = 1/15.
Can't find your answer?
Make a FREE account and ask your own questions, OR help others and earn volunteer hours!
Join our real-time social learning platform and learn together with your friends!
Join our real-time social learning platform and learn together with your friends!
Latest Questions
VERNEISH:
Two objects are moving along separate linear paths where each path is described by position, d, and time, t.
Maci:
what are the best stock apps for people who want to start investing?
Twaylor:
genuienly how do i get better self control i keep watcing youtube and theres this
IzzyCookie:
|dw:1792074056138:dw| Uhm, I don't know where to ask this, but I'm in college welding at my community college, and I need help learning vertical welding for
thehearken:
not sure if this is considered spam, im sorry if it is, but does anyone have tips
23 hours ago
1 Reply
0 Medals
5 days ago
3 Replies
2 Medals
1 week ago
3 Replies
0 Medals
3 days ago
9 Replies
1 Medal
1 week ago
2 Replies
1 Medal